# VideoAgent: Self-Improving Video Generation

### v2

@@ Abstract
Video generation has been used to generate visual plans for controlling robotic systems. Given an image observation and a language instruction, previous work has generated video plans which are then converted to robot controls to be executed. However, a major bottleneck in leveraging video generation for control lies in the quality of the generated videos, which often suffer from hallucinatory content and unrealistic physics, resulting in low task success when control actions are extracted from the generated videos. While scaling up dataset and model size provides a partial solution, integrating external feedback is both natural and essential for grounding video generation in the real world. With this observation, we propose VideoAgent for self-improving generated video plans based on external feedback. Instead of directly executing the generated video plan, VideoAgent first refines the generated video plans using a novel procedure which we call self-conditioning consistency, utilizing feedback from a pretrained vision-language model (VLM). As the refined video plan is being executed, VideoAgent collects additional data from the environment to further improve video plan generation. Experiments in simulated robotic manipulation from MetaWorld and iTHOR show that VideoAgent drastically reduces hallucination, thereby boosting success rate of downstream manipulation tasks. We further illustrate that VideoAgent can effectively refine real-robot videos, providing an early indicator that robotics can be an effective tool in grounding video generation in the physical world. Video demos can be found at https://video-as-agent.github.io.  [^1] Code available at https://github.com/Video-as-Agent/VideoAgent
@@ 1 Introduction
Large text-to-video models pretrained on internet-scale data have broad applications such as generating creative video content (Ho et al., 2022; Hong et al., 2022; Singer et al., 2022) and creating novel games (Bruce et al., 2024), animations (Wang et al., 2019), and movies (Zhu et al., 2023). Furthermore, recent work show that video generation can serve as simulators of the real-world (Yang et al., 2023b; Brooks et al., 2024), as well as policies with unified observation and action space (Du et al., 2024; Ko et al., 2023; Du et al., 2023). These recent applications of text-to-video generation models hold the great promise of internet-scale knowledge transfer (e.g., from generating human videos to generating robot videos), as well as paving the way to generalist agent (e.g., a single policy that can control multiple robots with different morphologies in different environments to perform diverse tasks).
@@
Nevertheless, text-to-video models have only had limited success in downstream applications in reality. For instance, in video generation as policy (Du et al., 2024; Ko et al., 2023), when an observation image and a language instruction are given to a video generation model, generated videos often hallucinate (e.g., objects randomly appear or disappear) or violate physical laws (e.g., a robot hand going through an object) (Yang et al., 2023b; Brooks et al., 2024). Such hallucinations and unrealistic physics have led to low task success rate when generated videos are converted to control actions through inverse dynamics models, goal conditioned policies, or other action extraction mechanisms (Wen et al., 2023; Yang et al., 2024; Ajay et al., 2024).
@@
While scaling up dataset and model size can be effective in reducing hallucination in large language models (LLMs) (Hoffmann et al., 2022), scaling is more difficult in video generation models. This is partially because language labels for videos are labor intensive to curate. Moreover, video generation has not converged to an architecture that is more favourable to scaling (Yang et al., 2024). Scaling aside, being able to incorporate external feedback to improve generation is one of the other most important breakthrough in LLMs (Ouyang et al., 2022b). It is therefore natural to wonder what kind of feedback is available for text-to-video models, and how we can incorporate these feedback to further improve the quality of the generated videos.
@@
To answer this question, we explore two types of feedback that are natural to acquire for video generation models, namely AI feedback from a vision-language model (VLM) and real-world execution feedback when generated videos are converted to motor controls. To utilize these feedback for self-improvement, we propose VideoAgent. Different from video generation as policy, which directly turns a generated video into control actions (Du et al., 2023; Ko et al., 2023), VideoAgent is trained to refine a generated video plan iteratively using feedback from a pretrained VLM. During inference, VideoAgent queries the VLM to select the best refined video plan, followed by execution of the plan in the environment. During online execution, VideoAgent observes whether the task was successfully completed and further improves the video generation model based on the execution feedback from the environment and additional data collected from the environment. The improvement to the generated video plan comes in two folds: First, we propose self-conditioning consistency for video diffusion model inspired by consistency models (Song et al., 2023; Heek et al., 2024), which enables low-quality samples from a video diffusion model to be further refined into high-quality samples. Second, when online access to the environment is available, VideoAgent executes the current video policy and collect additional successful trajectories to further finetune the video generation model on the successful trajectories. A visual illustration of VideoAgent is shown in Figure 1.
@@
We first evaluate the performance of VideoAgent in two simulated robotics manipulation environments, Meta-World (Yu et al., 2020) and iTHOR (Kolve et al., 2017), and show that VideoAgent improves task success across all environments and tasks evaluated. VideoAgent can even improve the success rate of difficult tasks by as much as 4X. Next, we provide a thorough study on the effect of different components in VideoAgent, including different ways to prompt the VLM and different types of feedback from the VLM, providing a recipe for utilizing VLM feedback for video generation. Lastly, we illustrate that VideoAgent can iteratively improve real-robot videos, providing early signal that robotics can be an important mean to ground video generation models in the real world.
@@ 2.2 Consistency Models
Diffusion models (Ho et al., 2020; Song et al., 2020b) have emerged as an important technique for data distribution modeling. During training, the model learns to map noisy data (at various noise levels) back to clean data in a single step. Concretely, let $x^{(0)}$ denote a clean image and $x^{(t)}$ denote the noisy image at noise level $t$, where $t\in[0,T]$, the training objective for a diffusion model $f_{\theta}(x^{(t)},t)$ can be written as
$\mathcal{L}_{\text{diffusion}}(\theta)=\mathbb{E}_{x^{(0)},\epsilon,t}\left[\|f_{\theta}(x^{(t)},t)-x^{(0)})\|^{2}\right],$ (2)
@@
$\mathcal{L}_{\text{consistency}}(\theta)=\mathbb{E}_{x^{(0)},\epsilon,t_{1},t_{2}}\left[\|f_{\theta}(x^{(t_{1})},t_{1})-\texttt{stopgrad}\big{(}{f_{\theta}(x^{(t_{2})},t_{2})}\big{)}\|^{2}\right],$ (3)
which encourages the output of the single-step map between different noise levels to be similar. In fact, both the diffusion loss in Equation 2 and the consistency loss in Equation 3 can be understood as exploiting the structure of the denoising procedure which corresponds to an ordinary differential equation (ODE). Specifically, as introduced in (Song et al., 2023; 2020a), the backward denoising procedure of a diffusion model can be characterized by an ODE, i.e.,
@@
with $s(x^{(t)},t)$ is some score function. During the entire path along $t\in(\epsilon,\infty]$, following this ODE should always maps $x_{t}$ to $x_{0}$. If we parametrize the model $f(x^{(t)},t)$ as the simulation following the ODE governed by $s(x^{(t)},t)$, we obtain the diffusion loss (2). Meanwhile, for all $t,t^{\prime}\in(\epsilon,\infty]$, we have $f(x^{(t)},t)=f(x^{(t^{\prime})},t^{\prime})$ along the simulation path, which induces the consistency loss (3). Therefore, we can combine the diffusion loss and the consistency loss together for model training, i.e.,
@@
3 Video Generation as Agent
@@
In this section, we introduce VideoAgent to improve video plan generation. In Section 3.1, we establish a new notion of consistency for video diffusion models, which we called self-conditioning consistency. In Section 3.2, we discuss how the video diffusion model trained with self-conditioning consistency can be used to refine generated video plans during inference. In Section 3.3, we discuss how VideoAgent further closes the self-improvement loop by collecting additional online data to further train the video generation and refinement model.
3.1 Video refinement through self-conditioning consistency
We first consider first-frame-and-language conditioned video generation following (Du et al., 2023; Ko et al., 2023), which finds a sequence of image frames for completing the task described by the language starting from the initial image. When a sample is drawn from a video generation model, it is commonly found that a part of the generated video (e.g., the beginning) is realistic while another part of the video (e.g., the end) hallucinates (Yang et al., 2023b). In other words, while a generated video plan may not fully complete the task specified, it provides meaningful information that can be further improved to achieve the correct plan. To leverage such partial progress, we consider a video consistency model to condition on previous self-generated samples to diffuse for the ground truth video, so that the model can learn to preserve the realistic part of the video while refining the hallucinatory part. Specifically, let $\mathbf{x}^{(0)}$ be a ground truth video, and $\hat{\mathbf{x}}$ be a generated video sample from the diffusion model. We define a self-conditioning consistency model as $\hat{f}_{\theta}(\hat{\mathbf{x}},\mathbf{x}^{(t)},t)$, which takes a generated video $\hat{\mathbf{x}}$ and a noisy version of the ground truth video $\mathbf{x}^{(t)}$ as input and predicts the clean video.
@@ 3.1 Video refinement through self-conditioning consistency
Figure 2: An illustration of Self-Conditioning Consistency model. Top row: denoising steps. Bottom row: refinement iterations. $\hat{\mathbf{x}}_{(i+1)}$ denotes the generated video plan at refinement iteration $(i+1)$. $\Delta$ denotes a fixed number of denoising steps. We condition the refinement on the generated video from the previous iteration $\hat{\mathbf{x}}_{(i)}$.
We make the observation that self-conditioning can be understood as a reparametrization of the implicit ODE solver for Equation 4  (Song et al., 2020a; Lu et al., 2022; Zhang & Chen, 2022; Chen et al., 2022). For example, Song et al. (2020a) considered the $1$st order ODE solver for Equation 4 following
$x^{(t-1)}=\sqrt{\alpha_{t-1}}\hat{x}+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\cdot s({x^{(t)},t}),$ (6)
where $\alpha_{t}$ follows the noise scheduling in Equation 2. Higher-order ODE solvers have also been considered in (Lu et al., 2022), all of which depend on the previous sample $\hat{x}$. Based on this observation, we introduce a previous video sample $\hat{\mathbf{x}}$ into our parametrization of $\hat{f}_{\theta}(\hat{\mathbf{x}},\mathbf{x}^{(t)},t)$ to mimic these ODE solvers. We emphasize that (Song et al., 2020a; Lu et al., 2022; Zhang & Chen, 2022) still follows the same parametrization for the score function $s({x^{(t)},t})$, which is learned by the vanilla diffusion loss. The self-conditioning ODE solver in Equation 6 is only exploited for accelerating generation. A visual illustration of self-conditioning is shown in Figure 2.
@@
We can learn the ODE solver through self-conditioning consistency by directly predicting the clean video $\mathbf{x}^{(0)}$ through a self-conditioning consistency
$\mathcal{L}_{\text{self-conditioning-consistency}}(\theta)$ $=\mathbb{E}_{\hat{\mathbf{x}},\mathbf{x}^{(0)},t}\left[\|\hat{f}_{\theta}(\hat{\mathbf{x}},\mathbf{x}^{(t)},t)-\mathbf{x}^{(0)}\|^{2}\right]$
$+\lambda\mathbb{E}_{\hat{\mathbf{x}}_{1},\hat{\mathbf{x}}_{2},t}\left[\|\hat{f}_{\theta}(\hat{\mathbf{x}}_{1},\mathbf{x}^{(t)},t)-\hat{f}_{\theta}(\hat{\mathbf{x}}_{2},\mathbf{x}^{(t)},t)\|^{2}\right],$ (7)
where $\hat{\mathbf{x}}_{1},\hat{\mathbf{x}}_{2}$ are two independent samples from the first-frame-and-language conditioned video generation model, and $\lambda$ is a hyperparameter for regularizing the similarity between different samples. To enable the “first guess” for $\hat{\mathbf{x}}$, we consider $f_{\theta}(\mathbf{x}^{(t)},t)$, which is still learned by the vanilla objective for video diffusion as
@@
$\mathcal{L}_{\text{video-diffusion}}(\theta)=\mathbb{E}_{\mathbf{x}^{(0)},\epsilon,t}\left[\|f_{\theta}(\mathbf{x}^{(t)},t)-\mathbf{x}^{(0)})\|^{2}\right].$ (8)
@@
The overall objective for training a self-conditioning-consistent video diffusion model boils down to
$\mathcal{L}(\theta)=\mathcal{L}_{\text{video-diffusion}}(\theta)+\lambda\mathcal{L}_{\text{self-conditioning-consistency}}(\theta).$ (9)
Note that despite the video generation model $f_{\theta}$ and the video refinement model $\hat{f}_{\theta}$ having different input arguments, we can nevertheless share the parameter between $f_{\theta}$ and $\hat{f}_{\theta}$ to train a single model for video generation and video refinement. We describe the training process for $f_{\theta}$ and $\hat{f}_{\theta}$ in Algorithm 1.
@@
While we can refine videos only from previously generated samples, it may be desirable to condition the refinement process on any additional feedback for the previously generated video that is available (e.g., feedback from humans or vision language models critiquing which part of the generated video is unrealistic). When such feedback is available, we can have the refinement model $\hat{f}$ further take the additional feedback as input to guide the refinement process, i.e.,
$\hat{f}_{\theta}(\mathbf{x},\mathbf{x}^{(t)},t,\text{feedback}),$ (10)
@@
After training the video generation model $f_{\theta}$ and the video refinement model $\hat{f}_{\theta}$ described in Equation 8 and Equation 7, we can sample from $f_{\theta}$ and iteratively apply $\hat{f}_{\theta}$ for video refinement. Specifically, let $\eta$ be the step size for the noise schedule, $\sigma_{t}$ be a time dependent noise term, VideoAgent first “guesses” the video plan using the first-frame-and-language conditioned video generation generation, i.e.,
$\mathbf{x}^{(t-1)}=\mathbf{x}^{(t)}-\eta\cdot\nabla_{\theta}f_{\theta}(\mathbf{x}^{(t)},t)+\sigma_{t}\cdot\epsilon$ (11)
@@
$\hat{\mathbf{x}}_{(i+1)}=\hat{f}_{\theta}(\hat{\mathbf{x}}_{(i)},\mathbf{x}^{(t)},t),$ (12)
@@
where $i$ denotes the video refinement iteration, with $\hat{\mathbf{x}}_{(0)}=\hat{\mathbf{x}}=\mathbf{x}^{(T)}$. We denote the final video after refinement as $\hat{\mathbf{x}}_{\text{refined}}$. A natural question is when to stop the iterative video refinement process. One option is to always refine for a fixed number of iterations. However, over-refinement may lead to less diverse output. To overcome this, we leverage a VLM as a proxy for the environment’s reward to assess whether a refined video is likely to lead to successful execution in the environment. Specifically, we denote a VLM as $\hat{\mathcal{R}}$, which takes a refined video $\hat{\mathbf{x}}_{(i)}$ and returns a binary value $\{0,1\}$ to determine whether a video is acceptable based on overall coherence, adherence to physical laws, and task completion (See prompt for VLM in Appendix A). With $\hat{\mathcal{R}}$, the refinement stops when the VLM decides that the refined video is acceptable. Namely, we have
@@
$\hat{\mathbf{x}}_{\text{refined}}=\hat{\mathbf{x}}_{(i^{*})},\quad\text{where}\quad i^{*}=\min\left\{i:\hat{\mathcal{R}}(\hat{\mathbf{x}}_{(i)})=1\right\}$ (13)
Algorithm 2 shows how video plans are generated, refined, and selected at inference time.
for iteration $=1$ to $N$ do
       Sample $\{(\mathbf{x}^{(0)},g)\}\sim\mathcal{D}$ and $t\sim\text{Uniform}(\{0,1,\dots,T\})$;
       Compute vanilla diffusion loss:
       $\mathcal{L}_{\text{video-diffusion}}=\left\|f_{\theta}(\mathbf{x}^{(t)},t)-\mathbf{x}^{(0)}\right\|^{2}$;
       Generate $\hat{\mathbf{x}}$ following Equation 11 and sample $\texttt{feedback}\sim\hat{\mathcal{R}}(\cdot|\hat{\mathbf{x}})$;
       Compute consistency loss:
       $\mathcal{L}_{\text{self-conditioning-consistency}}=\left\|\hat{f}_{\theta}(\hat{\mathbf{x}},\mathbf{x}^{(t)},t\,|\,\texttt{feedback})-\mathbf{x}^{(0)}\right\|^{2}$;
       Update parameters:
       $\theta\leftarrow\theta-\gamma\nabla_{\theta}\left(\mathcal{L}_{\text{video-diffusion}}+\mathcal{L}_{\text{self-conditioning-consistency}}\right)$;
3.3 Self-improvement through online finetuning
In addition to video refinement through self-conditioning consistency as described in Section 3.1, we can further characterize the combination of video generation and video refinement as a policy, which can be improved by training on additional real data collected from the environment during online interaction. Specifically, the goal is to maximize the expected returns of a policy through trial-and-error interaction with the environment:
where $\mathcal{R}$ is the true reward function, $\mathcal{E}$ is the interactive environment, and $\pi_{\theta}$ corresponds to Algorithm 2, which contains both the video generation model $f_{\theta}$ and the video refinement model $\hat{f}_{\theta}$ as learnable components to be improved.
where $\rho$ is the optical flow model that maps the refined video to low-level control actions. See Algorithm 3 for details of online policy finetuning.
[…]
       $\hat{\mathbf{x}}\leftarrow\pi_{\theta}(x_{0},g)$;
       for $i=0$ to max_refine_iterations do
             $\text{response}\leftarrow\hat{\mathcal{R}}(\hat{\mathbf{x}}_{(i)},g)$;
[…]
             $\hat{\mathbf{x}}_{(i+1)}\leftarrow\pi_{\theta}(\hat{\mathbf{x}}_{(i)},x_{0},g)$;
      $\text{success}\leftarrow\mathcal{R}(\rho(\hat{\mathbf{x}}_{\text{refined}}))$;
       if success then break;
[…]
Algorithm 2 VLM Guided Replan
for iteration $i=1$ to $N$ do
       $\mathcal{D}_{\text{new}}\leftarrow\emptyset$;
       for each $(\cdot,g)$ in $\mathcal{D}$ do
             $x_{0}\leftarrow\mathcal{E}.\text{reset}(g)$;
             $\hat{x}_{\text{refined}}\sim\pi_{\theta}(x_{0},g)$;
             if $\mathcal{R}(\rho(\hat{x}_{\text{refined}}))$ then
                   $\mathcal{D}_{\text{new}}\leftarrow\mathcal{D}_{\text{new}}\cup(\hat{x}_{\text{refined}},g)$;
      $\mathcal{D}\leftarrow\mathcal{D}\cup\mathcal{D}_{\text{new}}$;
       Finetune $\theta$ using Algorithm 1 on $\mathcal{D}$;
4 Experiments
We now evaluate the performance of VideoAgent, introducing the experimental settings and variants of VideoAgent in Section 4.1, end-to-end success rate of VideoAgent against the baselines in Section 4.2 and the effect of different components of VideoAgent in Section 4.3. Finally, we show that VideoAgent is effective in improving the quality of real robotic videos in Section 4.4.
4.1 Datasets and Experimental Setups
Datasets and Environments.
We follow the same evaluation setting as Ko et al. (2023), which considers three datasets: Meta-World (Yu et al., 2020), iTHOR (Kolve et al., 2017), and BridgeData V2 (Walke et al., 2023). Meta-World consists of 11 robotic manipulation tasks performed by a simulated Sawyer arm, with video demonstrations captured from three distinct camera angles. iTHOR is a simulated 2D object navigation benchmark, where an agent searches for specified objects across four room types. BridgeData V2 is a real-world dataset of robotic manipulation. See more details of datasets and environments in Appendix C.
Baselines and VideoAgent Variants.
We consider the following methods for comparison:
  • AVDC (baseline). This is the Actions from Video Dense Correspondences (AVDC) (Ko et al., 2023) baseline, which synthesizes a video and predicts optical flow to infer actions.
  • VideoAgent-Online-Replan. This variant incorporates online filtering of successful trajectories with the replanning mechanism, where replanning is conducted first, and more successful trajectories after replanning are added back to the training data.
4.2 End-to-End Task Success

### v3

@@ Abstract
Video generation has been used to generate visual plans for controlling robotic systems. Given an image observation and a language instruction, previous work has generated video plans which are then converted to robot controls to be executed. However, a major bottleneck in leveraging video generation for control lies in the quality of the generated videos, which often suffer from hallucinatory content and unrealistic physics, resulting in low task success when control actions are extracted from the generated videos. While scaling up dataset and model size provides a partial solution, integrating external feedback is both natural and essential for grounding video generation in the real world. With this observation, we propose VideoAgent for self-improving generated video plans based on external feedback. Instead of directly executing the generated video plan, VideoAgent first refines the generated video plans using a novel procedure which we call self-conditioning consistency, allowing inference-time compute to be turned into better generated video plans. As the refined video plan is being executed, VideoAgent can collect additional data from the environment to further improve video plan generation. Experiments in simulated robotic manipulation from MetaWorld and iTHOR show that VideoAgent drastically reduces hallucination, thereby boosting success rate of downstream manipulation tasks. We further illustrate that VideoAgent can effectively refine real-robot videos, providing an early indicator that robots can be an effective tool in grounding video generation in the physical world. Video demos and code can be found at https://video-as-agent.github.io.
Machine Learning, ICML
@@ 1 Introduction
Large text-to-video models pretrained on internet-scale data have broad applications such as generating creative video content (Ho et al., 2022; Hong et al., 2022; Singer et al., 2022) and creating novel games (Bruce et al., 2024), animations (Wang et al., 2019), and movies (Zhu et al., 2023). Furthermore, recent work show that video generation can serve as simulators of the real-world (Yang et al., 2023; Brooks et al., 2024), as well as policies with unified observation and action space (Du et al., 2024; Ko et al., 2023; Du et al., 2023). These recent applications of text-to-video generation models holds great promise of internet-scale knowledge transfer (e.g., from generating human videos to generating robot videos), as well as paving the way to generalist agent (e.g., a single policy that can control multiple robots with different morphologies in different environments to perform diverse tasks).
@@
Nevertheless, text-to-video models have only had limited success in downstream applications in reality. For instance, in video generation as policy (Du et al., 2024; Ko et al., 2023), when an observation image and a language instruction are given to a video generation model, generated videos often hallucinate (e.g., objects randomly appear or disappear) or violate physical laws (e.g., a robot hand going through an object) (Yang et al., 2023; Brooks et al., 2024). Such hallucinations and unrealistic physics have led to low task success rate when generated videos are converted to control actions through inverse dynamics models, goal conditioned policies, or other action extraction mechanisms (Wen et al., 2023; Yang et al., 2024; Ajay et al., 2024).
@@
While scaling up dataset and model size can be effective in reducing hallucination in large language models (LLMs) (Hoffmann et al., 2022), scaling is more difficult in video generation models. This is partially because language labels for videos are labor intensive to curate. Moreover, video generation has not converged to an architecture that is more favourable to scaling (Yang et al., 2024). Scaling aside, being able to incorporate external feedback to improve generation is one of the other most important breakthrough in LLMs (Ouyang et al., 2022). It is therefore natural to wonder what kind of feedback is available for text-to-video models, and how we can incorporate these feedback to further improve the quality of the generated videos.
@@
To answer this question, we explore two types of feedback that are natural to acquire for video generation models, namely AI feedback from a vision-language model (VLM) and real-world execution feedback when generated videos are converted to motor controls. To utilize these feedback for self-improvement, we propose VideoAgent. Different from video generation as policy, which directly turns a generated video into control actions (Du et al., 2023; Ko et al., 2023), VideoAgent is trained to refine a generated video plan iteratively using feedback from a pretrained VLM. During inference, VideoAgent queries the VLM to select the best refined video plan, allowing inference-time compute to be turned into better generated video plans, followed by execution of the plan in the environment. During online execution, VideoAgent observes whether the task was successfully completed, and further improves the video generation model based on the execution feedback from the environment and additional data collected from the environment. The improvement to the generated video plan comes in three folds: First, we propose self-conditioning consistency for video diffusion model inspired by consistency models (Song et al., 2023; Heek et al., 2024), which enables low-quality samples from a video diffusion model to be further refined into high-quality samples. Second, VLM feedback combined with more inference-time compute leads to better video plans. Lastly, when online access to the environment is available, VideoAgent executes the current video plan and collects additional successful trajectories to further finetune the video generation model. A visual illustration of VideoAgent is shown in Figure 1.
@@
We first evaluate VideoAgent in two simulated robotic manipulation environments, Meta-World (Yu et al., 2020) and iTHOR (Kolve et al., 2017), and show that VideoAgent improves task success across all environments and tasks evaluated. Next, we provide a thorough study on the effect of different components in VideoAgent, including different types of feedback from the VLM, providing a recipe for utilizing VLM feedback for video generation. Lastly, we illustrate that VideoAgent can iteratively improve real-robot videos, providing early signal that robotics can be an important mean to ground video generation models in the real world.
@@ 2.2 Consistency Models
Diffusion models (Ho et al., 2020; Song et al., 2020b) have emerged as an important technique for generative modeling of high-dimensional data. During training, a diffusion model learns to map noisy data (at various noise levels) back to clean data in a single step. Concretely, let $x^{(0)}$ denote a clean image and $x^{(t)}$ denote the noisy image at noise level $t$, where $t\in[0,T]$, the training objective for a diffusion model $f_{\theta}(x^{(t)},t)$ can be written as
$\mathcal{L}_{\text{diffusion}}(\theta)=\mathbb{E}_{x^{(0)},\epsilon,t}\left[\|f_{\theta}(x^{(t)},t)-x^{(0)}\|^{2}\right],$ (2)
@@
$\mathcal{L}_{\text{consistency}}(\theta)$ $=\mathbb{E}_{x^{(0)},\epsilon,t_{1},t_{2}}\Big{[}\|f_{\theta}(x^{(t_{1})},t_{1})$
$\quad-\texttt{stopgrad}\big{(}f_{\theta}(x^{(t_{2})},t_{2})\big{)}\|^{2}\Big{]},$ (3)
which encourages the output of the single-step map between different noise levels to be similar. In fact, both the diffusion loss in Equation 2 and the consistency loss in Equation 3 can be understood as exploiting the structure of the denoising procedure which corresponds to an ordinary differential equation (ODE). Specifically, as introduced in (Song et al., 2023, 2020a), the backward denoising procedure of a diffusion model can be characterized by an ODE, i.e.,
@@
with $s(x^{(t)},t)$ is some score function. During the entire path along $t\in(\epsilon,\infty]$, following this ODE should always map $x^{(t)}$ to $x^{(0)}$. If we parametrize the model $f(x^{(t)},t)$ as the simulation following the ODE governed by $s(x^{(t)},t)$, we obtain the diffusion loss (2). Meanwhile, for all $t,t^{\prime}\in(\epsilon,\infty]$, we have $f(x^{(t)},t)=f(x^{(t^{\prime})},t^{\prime})$ along the simulation path, which induces the consistency loss (3). Therefore, we can combine the diffusion loss and the consistency loss together for model training, i.e.,
@@
3 Video Generation as an Agent
@@
In this section, we introduce VideoAgent, a framework for improving video plan generation. In Section 3.1, we develop self-conditioning consistency to iteratively refine generated video plans. In Section 3.2, we describe how a diffusion model trained with self-conditioning consistency can leverage inference-time compute to refine generated video plans. Finally, in Section 3.3, we illustrate how VideoAgent closes the self-improvement loop by collecting additional online data to further enhance video generation and refinement.
3.1 Refinement through self-conditioning consistency
We consider first-frame-and-language conditioned video generation following (Du et al., 2023; Ko et al., 2023), which generates a sequence of image frames to complete the task described by the language starting from the initial image. In practice, generated videos often contain hallucinations  (Yang et al., 2023). While such inaccuracies may prevent a video plan from fully completing the task, the generated video may still make meaningful progress towards completing the task. Therefore, instead of independently sampling many videos and hoping that one may be free from hallucinations, we propose to refine previously generated videos iteratively.
Specifically, let $\mathbf{x}^{(0)}$ denote the ground truth video and $\hat{\mathbf{x}}$ a generated video from the original text-to-video diffusion model. We introduce a self-conditioning consistency model, $\hat{f}_{\theta}(\hat{\mathbf{x}},\mathbf{x}^{(t)},t)$, which takes a generated video $\hat{\mathbf{x}}$ and a noisy version of the ground truth $\mathbf{x}^{(t)}$ as inputs to predict the clean video. This formulation enables iterative refinement by conditioning the model on its previous predictions, as illustrated in Figure 2. We denote video samples from the refinement model after the $i$-th iteration as $\hat{\mathbf{x}}_{i}$. Self-conditioning is inspired by a reparameterization of the implicit ODE solver for Equation 4  (Song et al., 2020a; Lu et al., 2022; Zhang & Chen, 2022; Chen et al., 2022). For instance, (Song et al., 2020a) considered the first-order ODE solver for Equation 4 following:
$\mathbf{x}^{(t-1)}=\sqrt{\alpha_{t-1}}\mathbf{x^{(0)}}+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\cdot s({\mathbf{x}^{(t)},t}).$ (6)
@@ 3.1 Refinement through self-conditioning consistency
Figure 2: An illustration of Self-Conditioning Consistency. The horizontal direction represents the regular denoising process. The two rows represent two refinement iterations. $\hat{\mathbf{x}}_{i}$ denotes the generated video plan at refinement iteration $i$. We condition the refinement iteration $i+1$ on the generated video from the previous iteration $\hat{\mathbf{x}}_{i}$.
In VideoAgent, we adapt Equation 6 by replacing the $\mathbf{x}^{(0)}$ term with $\mathbf{\hat{x}}$, the previously generated video sample, as illustrated in Figure 2. In standard DDIM-based methods (Song et al., 2020a), $\mathbf{x}^{(0)}$ is typically obtained as an intermediate estimate from $\mathbf{x}^{(t)}$ within the same iteration. In contrast, our approach reuses $\mathbf{\hat{x}}$ from a previous iteration, allowing for a self-conditioning mechanism that improves temporal coherence. By enforcing consistency across iterations, our method enables the denoising process to correct potential failures more effectively.
@@
We learn the ODE solver through self-conditioning consistency by directly predicting the clean video $\mathbf{\hat{x}_{i+1}}$ using:
$\mathcal{L}_{\text{SC-consistency}}(\theta)$ $=\mathbb{E}_{\hat{\mathbf{x}},\mathbf{x}^{(0)},t}\Big{[}\|\hat{f}_{\theta}(\hat{\mathbf{x}},\mathbf{x}^{(t)},t)-\mathbf{x}^{(0)}\|^{2}\Big{]}$
$\quad+\mu\mathbb{E}_{\hat{\mathbf{x}}_{1},\hat{\mathbf{x}}_{2},t}\Big{[}\|\hat{f}_{\theta}(\hat{\mathbf{x}}_{1},\mathbf{x}^{(t)},t)$
$\quad-\hat{f}_{\theta}(\hat{\mathbf{x}}_{2},\mathbf{x}^{(t)},t)\|^{2}\Big{]}.$ (7)
The first term in Equation 7 represents the standard diffusion loss with the additional conditioning on $\hat{\mathbf{x}}$, while the second term regularizes the similarity between different refinement iterations ($\mathbf{\hat{x}}_{1}$ and $\mathbf{\hat{x}}_{2}$) to promote coherence across iterations. This iterative refinement process distinguishes self-conditioning consistency from traditional consistency models. Combined with the standard objective for video diffusion:
@@
$\mathcal{L}_{\text{video-diffusion}}(\theta)=\mathbb{E}_{\mathbf{x}^{(0)},\epsilon,t}\left[\|f_{\theta}(\mathbf{x}^{(t)},t)-\mathbf{x}^{(0)})\|^{2}\right],$ (8)
@@
the overall objective for training a self-conditioning-consistent video diffusion model thus becomes:
$\mathcal{L}(\theta)=\mathcal{L}_{\text{video-diffusion}}(\theta)+\lambda\mathcal{L}_{\text{SC-consistency}}(\theta).$ (9)
Note that while the video generation model $f_{\theta}$ and the video refinement model $\hat{f}_{\theta}$ have different input arguments (first frame versus previously generated video), we can share their parameters to train a single unified model for both video generation and refinement tasks. This parameter-sharing approach allows us to leverage the same model architecture for generating initial video plans and iterative refinement. The training process for $f_{\theta}$ and $\hat{f}_{\theta}$ is detailed in Algorithm 1 in Appendix A.
@@
While we can refine videos only from previously generated samples, it may be desirable to condition the refinement process on any additional feedback for the previously generated video that is available (e.g., feedback from humans or vision language models critiquing which part of the generated video is unrealistic). When such feedback is available, we can have the refinement model $\hat{f}$ further take the additional feedback as input, combined with the task description, to guide the refinement process, i.e.,
$\hat{f}_{\theta}(\mathbf{x},\mathbf{x}^{(t)},t|\text{feedback}),$ (10)
@@
After training the video generation model $f_{\theta}$ and the video refinement model $\hat{f}_{\theta}$ described in Equation 8 and Equation 7, we can sample from $f_{\theta}$ and iteratively apply $\hat{f}_{\theta}$ for video refinement. Specifically, let $\eta$ be the step size for the noise schedule, $\sigma_{t}$ be a time dependent noise term, VideoAgent first generates a video plan through
$\mathbf{x}^{(t-1)}=\mathbf{x}^{(t)}-\eta\cdot\nabla_{\theta}f_{\theta}(\mathbf{x}^{(t)},t)+\sigma_{t}\cdot\epsilon.$ (11)
@@
$\hat{\mathbf{x}}_{i+1}=\hat{f}_{\theta}(\hat{\mathbf{x}}_{i},\mathbf{x}^{(t)},t),$ (12)
@@
where $i$ denotes the video refinement iteration, with $\hat{\mathbf{x}}_{0}=\hat{\mathbf{x}}=\mathbf{x}^{(T)}$. We denote the final video after refinement as $\hat{\mathbf{x}}_{\text{refined}}$. A natural question is when to stop the iterative video refinement process. We use a VLM as a proxy for the environment’s reward to assess whether a refined video is likely to lead to successful execution in the environment. Specifically, we denote a VLM as $\hat{\mathcal{R}}$, which takes a refined video $\hat{\mathbf{x}}_{i}$ and returns a binary value $\{0,1\}$ to determine whether a video is acceptable based on overall coherence, adherence to physical laws, and task completion (See prompt for VLM in Appendix B). With $\hat{\mathcal{R}}$, the refinement stops when the VLM decides that the refined video is acceptable. Namely, we have
@@
$\hat{\mathbf{x}}_{\text{refined}}=\hat{\mathbf{x}}_{i^{*}},\quad\text{where}\quad i^{*}=\min\left\{i:\hat{\mathcal{R}}(\hat{\mathbf{x}}_{i})=1\right\}$ (13)
Algorithm 2 in Appendix A shows details for video plan generation, refinement, and selection during inference.